Part A Review A proton is released from rest in a uniform electric field of magnitude 2.18x105 N/C. Find the speed of the proton after it has traveled (a) 1.10 cm and (b) 14.0 cm. m/s Submit Request Answer Part B m/s
Answer #1

Proton is released from rest so u = 0 m/s = initial speed

E = 2.18 x 105 N/C

F = E*q
F = 2.18x105*(1.6x10-19)
F = 3.488 x 10-14 N

and we also know that

F=m*a
a = F/m
a = (3.488 x 10-14)/(1.672 x 10-27)
a = 2.086124 x 1013 m/s2

using a in

v2 = u2 + 2*a*s

v = (2*a*s)

v = (2*2.086124 x 1013*0.011)

v = 677456.54 m/s

v = 6.77 x 105 m/s

(b)

same for this part s = 0.14 m

v2 = u2 + 2*a*s

v = (2*a*s)

v = (2*2.086124 x 1013*0.14)

v = 2416846.77 m/s

v = 2.41 x 106 m/s